# The interplay of the parametric instability and MRI in oscillatory shear flows

### v1

@@ Abstract
The evolution of warped disks is governed by internal, oscillatory shear flows driven by their distorted geometry. However, these flows are known to be vigorously unstable to a hydrodynamic parametric instability. In many warped systems this might coexist and compete with the magnetorotational instability. The interplay of these phenomena and their combined impact on the internal flows has not been studied. To this end we perform three-dimensional, magneto+hydrodynamic unstratified shearing box simulations with an oscillatory radial forcing function, to mimic the effects of a warped disk. In the hydrodynamic study we find that the parametric instability manifests as strong, vertical ‘elevator’ flows which resist the sloshing motion. Above a critical forcing amplitude, these also emerge in our magnetized runs and dominate the vertical stress, although they are partially weakened by the MRI and hence the system equilibrates with larger radial sloshing flows. Below this critical forcing, the MRI effectively quenches the parametric instability. In all cases we find that the internal stresses are anisotropic in character and better described by a viscoelastic relationship with the shearing flows. Unfortunately, these important effects are typically unresolved in global simulations of warped disks and are simplified in analytically tractable models. The incorporation of such complex, warp-amplitude-dependent, viscoelastic stresses will sensitively regulate the laminar flow response, and inevitably modify the detailed spatio-temporal evolution of warped systems.
@@ 1 Introduction
Distorted astrophysical disks, which depart from the coplanar and circular assumptions of standard models, are now thought to be ubiquitous. Indeed, whenever there is some inherent misalignment present in a system, this can cause the disk to become warped, wherein the orbital plane varies as a function of radius. For example, this misalignment might derive from the Lense-Thirring torque around tilted spinning black holes (Bardeen & Petterson, 1975), the inclined delivery of material during infall events (Bate et al., 2010; Kuffmeier et al., 2021), the gravitational perturbations from tilted companions or planets (Papaloizou & Terquem, 1995; Xiang-Gruess & Papaloizou, 2013; Facchini et al., 2013) or even inclined magnetic dipoles interacting with the surrounding disk (Lai, 1999; Romanova et al., 2021). This diversity of physical mechanisms is borne out in the varied observational evidence for warps. Warps have been indirectly inferred from the superorbital modulation of X-ray binary light sources (Katz, 1973; Kotze & Charles, 2012) and the shadows cast onto the outer regions of protoplanetary disks by inner warped regions (Marino et al., 2015; Casassus et al., 2018). Furthermore, warps have even been invoked to explain the variability of active galactic nuclei (AGN) (Raj & Nixon, 2021; Kaaz et al., 2023) and the precession of jets (Miller-Jones et al., 2019; Cui et al., 2023). Tantalizing direct evidence for warps has also been provided by the warped maser emission in NGC 4258 (Miyoshi et al., 1995) and the distorted midplane of the protostellar disks L1489 IRS (Sai et al., 2020) and IRAS 04368+2557 (Sakai et al., 2019).
@@
This prevalence for warped systems demands a rigorous theoretical understanding to model their evolution in space and time. Whilst the early studies laid much of the geometrical framework for global warp evolution (see Petterson, 1977a, b, 1978), they assumed that the disk behaves as a series of rigid, concentric annuli which interact in a purely viscous fashion. The paradigm was fundamentally changed by Papaloizou & Pringle (1983) who realized that the distorted geometry actually establishes pressure gradients which drive internal, radial oscillatory shear flows that are odd about the midplane and efficiently communicate angular momentum. The pressure forces experienced by a fluid parcel vary over the orbital timescale as it traverses the warped geometry. This variation yields a near-resonant forcing, which matches the natural epicylic frequency. To regulate the resonant response, analytical work typically invokes the viscous $\alpha$ prescription (Shakura & Sunyaev, 1973) as a source of damping. Indeed, provided $\alpha$ exceeds the aspect ratio of the disk $H/R$, the viscosity effectively moderates the internal flows and sets up a quasi-steady ‘sloshing’ state whilst the warp evolves over longer timescales – this is the so-called diffusive regime. In the opposite case that $\alpha<H/R$, the weaker viscosity is unable to temper the internal flows and instead the warp is detuned from resonance by dynamically evolving as a bending wave in the wave-like regime (Papaloizou & Lin, 1995). These linear theories were subsequently extended towards nonlinear warp amplitudes by Ogilvie (1999) for the diffusive regime and by Ogilvie (2006, 2018) for the wave-like regime. Efforts to combine both regimes into a generalized set of evolutionary equations have also recently been pursued (Martin et al., 2019; Dullemond et al., 2022).
@@
All of these theories rely on the simplification that the disk flow structure is laminar and smooth and that any small-scale turbulence is well parameterized by the viscous $\alpha$ model. With this in mind, a number of global, laminar numerical experiments have found good agreement with the theoretical equations for warp evolution in both regimes (e.g. Lodato & Price, 2010; Fragner & Nelson, 2010; Facchini et al., 2013; Nealon et al., 2015; Kimmig & Dullemond, 2024; Fairbairn & Ogilvie, 2023). However, such a simple viscous closure model might not capture the true behavior of the underlying turbulence and its interplay with the warp. Indeed, it has long been appreciated that distorted disks are susceptible to a hydrodynamic, parametric instability which feeds off the internal, oscillatory shear flows (Papaloizou & Terquem, 1995; Gammie et al., 2000). This small scale instability is most readily understood within a local warped shearing box model (Ogilvie & Latter, 2013a, b) which was numerically implemented in the two-dimensional simulations of Paardekooper & Ogilvie (2019). They recovered the linear growth phase of the instability and the ensuing nonlinear saturated state. This significantly modified the laminar flows and reduced the associated angular momentum transport in the warped disk. Furthermore, they found complicated variability in the final turbulent state, as a function of warp amplitude, and also evidence of hysteresis as the saturated flows depended on the warp history. This instability can also efficiently extract energy from the sloshing flows and lead to a rapid damping of the driving warp (Fairbairn & Ogilvie, 2023). Unfortunately, this richness of dynamics is typically unresolved in global numerical experiments. The only clear identification so far has been the high resolution, 120 million Lagrangian particle runs of Deng et al. (2020); Deng & Ogilvie (2022) which attain $\sim 8$ elements per scale height. The emergence of the parametric instability dramatically damps the warp, strongly departing from the laminar prescriptions.
@@
This purely hydrodynamic picture is further complicated by the possible presence of magnetic fields in warped disks – relevant to the case of X-ray binaries or AGN disks for example. Analytical progress on this front is limited to the regime of strong magnetization for which vertical field lines threading the disk provide an effective tension – resisting the oscillatory, shearing flows and detuning the aforementioned Keplerian resonance (Paris & Ogilvie, 2018). In the alternate case of weak fields the disk is readily unstable to the magnetorotational instability (MRI) (Balbus & Hawley, 1991; Hawley & Balbus, 1991). Whilst this mechanism is often invoked to explain radial transport of mass and angular momentum, the turbulence might also help moderate the vertically shearing, sloshing flows. Indeed, often the in-plane accretion stresses are equated to some value of effective $\alpha$, which is then used in the isotropic, viscous theories outlined above. In reality the picture is inevitably more complicated as the MRI turbulence is anisotropic and perhaps doesn’t respond instantaneously to the local shear rate. Therefore we need to appeal to numerical simulations to gain further insight. Indeed, the global, magnetized warped torus simulations of Sorathia et al. (2013a) suggested a distinct evolution from that predicted by viscous theory. Furthermore, they found that the flow was dominated by hydrodynamic rather than magnetized stresses, akin to their purely hydro runs in Sorathia et al. (2013b). Contrastingly, Nealon et al. (2016) were able to reproduce the MHD simulations of a warped disk around a spinning black hole (Krolik & Hawley, 2015) using viscous simulations. Despite these global efforts, there has been little work to investigate in detail how the local internal flows are modified by the MRI. Exceptionally, Torkelsson et al. (2000) performed three-dimensional stratified, shearing box simulations of an MRI turbulent disk where the effects of the warp are mocked up via an imposed epicylic, sloshing initial condition that is odd about the midplane. This controlled experiment identified the early emergence of the parametric instability, followed by exponential damping of the sloshing motions due primarily to the hydrodynamic transport of momentum. The damping timescales found are consistent with vertical viscosities which are (coincidentally) comparable to the in-plane MRI effective viscosity – thus supporting an isotropic $\alpha$.
@@
Despite this progress, there are several routes demanding further inquiry. The MRI simulations of Torkelsson et al. (2000) implement the warp as a free initial condition which quickly decays. Therefore, there is little time to establish a saturated, magnetized turbulent state in which the parametric instability, MRI and sloshing flows come into an equilibrium. Indeed, extracting the nature of the turbulent stresses is complicated by the limited timescales. Contrastingly, the detailed parametric instability study of Paardekooper & Ogilvie (2019) establishes steady sloshing states but focuses on two-dimensional hydrodynamics. In this paper we will find a compromise between both approaches. Specifically, we will compare controlled hydrodynamic and magnetized, local simulations of oscillatory shear flows which are maintained by a time-dependent forcing function, mimicking the radial pressure gradients set up by warped disks. In this initial study we will further simplify matters by considering the unstratified problem, which will be extended towards the stratified case in future work. The outline for this paper is as follows. In section 2 we will describe the problem setup and summarize the laminar model predictions. In section 3 we will benchmark our numerical setup by capturing the linear growth of the parametric instability and compare with theoretical predictions. In section 4 we will examine the fiducial hydrodynamic sloshing run and study the emerging stresses. This will be compared with the fiducial magnetized sloshing run in section 5, before we perform a study varying the forcing amplitude. We will discuss our findings in the context of previous studies in section 6 and finally conclude in section 7.
@@ 2 Problem Setup
where $\rho$ is the density, $\mathbf{u}$ is the velocity vector and $\mathbf{B}$ is the magnetic field. The total pressure $P^{*}=p+\mathbf{B}\cdot\mathbf{B}/2$, is composed of the magnetic pressure and the thermal pressure p. Throughout this work, we will assume the simplest isothermal equation of state (EoS) $p=c^{2}\rho$, connecting pressure and density via the sound speed $c$. This removes the need to explicitly solve an energy equation. On the right-hand side of equation (2) we see contributions from the Coriolis force and the tidal potential, featuring the shear parameter $q=-d\ln\Omega/d\ln r=3/2$, for a Keplerian disk. Furthermore, the shear viscous stress tensor (Landau & Lifshitz, 1959) is defined as
@@
$\mathsf{T}_{ij}=\rho\nu_{ij}\left(\frac{\partial u_{i}}{\partial x_{j}}-\frac{\partial u_{j}}{\partial x_{i}}-\frac{2}{3}\delta_{ij}\nabla\cdot\mathbf{u}\right),$ (4)
@@
where we allow for an anisotropic kinematic viscosity $\nu_{ij}$. [^1] Note that this simple prescription for characterizing anisotropy means that $\mathsf{T}_{ij}$ does not generally transform in a consistent tensorial fashion. Another non-ideal effect is included in the induction equation (3) when the ohmic resistivity $\eta$ is set to be non-zero. Note that for the purposes of this focused study, we have neglected the vertical component of gravity and therefore consider unstratified disk setups. This extension will be addressed in future work.
@@
Figure 1: Global and local cartoon of warped disk. Left: the global warped disk consists of a series of orbits with a radially varying orientation of angular momentum. Right: a zoomed-in patch of the warped disk manifests as a tilting of the midplane. This projects a component of the vertical pressure gradient in the radial direction and drives sloshing flows.
@@
This study is primarily interested in probing the effect of instabilities and ensuing nonlinear turbulence in moderating oscillatory flows forced by distorted geometries. In particular, it has long been appreciated that warped disk evolution is driven by self-induced internal flows (e.g. Papaloizou & Pringle, 1983; Ogilvie & Latter, 2013a). To gain some heuristic intuition, consider the warped disk cartoon in Fig. 1. The globally distorted geometry is visualised on the left where the midplane changes inclination as a function of radius. The orientation of the angular momentum on each disk annulus is formally encoded by the radial and temporal variation of the unit tilt vector $\hat{\mathbf{l}}(r,t)$. The warp amplitude is then quantified by $\psi=\left|d\hat{\mathbf{l}}/d\ln r\right|$. On the right, we zoom in to a portion of the disk and examine the local manifestation of the warp as a tilted midplane, defined by $z_{\rm mid}=\psi x_{\rm mid}$. In a flat, stratified disk, the vertical hydrostatic pressure gradient is given by $\partial p/\partial z=-\rho\Omega_{0}^{2}z$. Therefore, upon tilting the midplane, a portion of this force is projected into the radial direction, yielding $-\mathbf{\hat{x}}\cdot\nabla p\sim\rho\Omega_{0}^{2}\psi z$ (Lodato & Pringle, 2007). As we track the orbit around the disk, the midplane tilts back and forth, incurring a sinusoidal dependence in the radial forcing such that it has the form $f_{\rm w}\sim\rho\Omega_{0}^{2}\psi z\cos(\Omega_{0}t)$. The characteristic properties of this forcing is that it is linear in the vertical coordinate $z$ and oscillates on the orbital timescale.
@@
Of course, the simple, unstratified shearing box setup introduced above precludes the self-consistent generation of these pressure gradients. Instead, we will implement an artificial forcing function that emulates key properties in a controlled fashion. To this end we will add
@@
to the right hand side of equation (2) where $\psi_{\rm f}$ is now a complex number allowing for an arbitrary choice of phase. Here, $L_{z}$ is the vertical domain in our shearing box, whilst $H\equiv c/\Omega_{0}$ sets the characteristic scale-height. Near the box midplane, this form approximates a linear function of $z$ before turning around to satisfy periodic boundary conditions. This idealized model, although simple, contains all of the fundamental physics for driving sloshing flows in warped disks. An alternative approach might be to use the warped generalization of the shearing box, developed by Ogilvie & Latter (2013a), which explicitly incorporates the warped geometry via Lagrangian coordinates which track the tilted orbital motion. However, in order to leverage current state-of-the-art MHD numerical code bases, with inbuilt support for standard shearing boxes, we will adopt our simplified approach.
To gain some intuition about the growth of sloshing motions in our setup, consider the hydrodynamic limit of equations (1) and (2), equipped with the forcing function (5). Furthermore, we will assume that the flow structure is horizontally homogeneous, laminar, and moderated by the shear viscosity, which acts as a model for the turbulent stresses investigated later in this paper. Motivated by the sinusoidal vertical structure of $\mathbf{f}_{\rm w}$ the velocity response follows
@@
where $u_{\rm K}=-q\Omega_{0}x$ is the usual Keplerian shear flow solution in the absence of warped forcing and we have defined $k_{\rm b}\equiv 2\pi/L_{z}$. It is also sometimes useful to work with the velocity relative to the background Keplerian flow which we will denote, $\mathbf{v}=\mathbf{u}-\mathbf{u}_{\rm k}$. Inserting this ansatz into the governing equations yields
@@
In the simplest case for which the system is unforced and inviscid, the dynamics reduces to the epicylic oscillator solution,
@@
for which $S$ sets the sloshing amplitude and phase. The vertical layers oscillate independently as a series of stacked epicylic orbits. The amplitude of these oscillations as a function of $z$ is enforced by our ansatz in equations (6) and (7), although for this free sloshing, any arbitrary function of $z$ is obviously permitted. We will restrict ourselves to this coherent sinusoidal structure which is established by our forcing function. A useful diagnostic which encodes the magnitude of this $x$-$y$ sloshing motion is the epicylic amplitude
@@
When the system is inviscid but now forced, equation (2) reduces to the standard, resonantly forced oscillator with a linearly growing response,
@@
This resonance between the warp-induced radial forcing frequency and the orbital frequency (equal to the natural epicyclic frequency for our Keplerian shear), underlies the difficulty in developing a theory for warps in the low viscosity, bending wave regime, where the response is dynamical and no quasi-steady sloshing state can be found. Meanwhile, the inclusion of viscosity tempers the resonance and gives the steady state sloshing
@@
$B_{\rm vis}(t)=-\frac{-\Omega_{0}^{2}\psi_{\rm f}}{2\left[\nu_{xz}\nu_{yz}k_{\rm b}^{4}+i\Omega_{0}(\nu_{xz}+\nu_{yz})k_{\rm b}^{2}\right]}e^{i\Omega_{0}t}.$ (16)

### v2

# The Interplay of Parametric and Magnetorotational Instabilities in Oscillatory Shear Flows

@@ Abstract
The evolution of warped disks is governed by internal, oscillatory shear flows driven by their distorted geometry. However, these flows are known to be vigorously unstable to a hydrodynamic parametric instability. In many warped systems, this might coexist and compete with the magnetorotational instability (MRI). The interplay of these phenomena and their combined impact on the internal flows has not been studied. To this end, we perform three-dimensional, magnetohydrodynamic unstratified shearing box simulations with an oscillatory radial forcing function to mimic the effects of a warped disk. In the hydrodynamic study, we find that the parametric instability manifests as strong, vertical ‘elevator’ flows that resist the sloshing motion. Above a critical forcing amplitude, these also emerge in our magnetized runs and dominate the vertical stress, although they are partially weakened by the MRI, and hence the system equilibrates with larger radial sloshing flows. Below this critical forcing, the MRI effectively quenches the parametric instability. In all cases, we find that the internal stresses are anisotropic in character and better described by a viscoelastic relationship with the shearing flows. Unfortunately, these important effects are typically unresolved in global simulations of warped disks and are simplified in analytically tractable models. The incorporation of such complex, warp-amplitude-dependent, viscoelastic stresses will sensitively regulate the laminar flow response and inevitably modify the detailed spatio-temporal evolution of warped systems.
@@ 1 Introduction
Distorted astrophysical disks, which depart from the coplanar and circular assumptions of standard models, are now thought to be ubiquitous. Indeed, whenever there is some inherent misalignment present in a system, this can cause the disk to become warped, wherein the orbital plane varies as a function of the radius. For example, this misalignment might derive from the Lense-Thirring torque around tilted spinning black holes (Bardeen & Petterson, 1975), the inclined delivery of material during infall events (Bate et al., 2010; Kuffmeier et al., 2021), the gravitational perturbations from tilted companions or planets (Papaloizou & Terquem, 1995; Xiang-Gruess & Papaloizou, 2013; Facchini et al., 2013) or even inclined magnetic dipoles interacting with the surrounding disk (Lai, 1999; Romanova et al., 2021). This diversity of physical mechanisms is borne out in the varied observational evidence for warps. Warps have been indirectly inferred from the superorbital modulation of X-ray binary light sources (Katz, 1973; Kotze & Charles, 2012) and the shadows cast onto the outer regions of protoplanetary disks by inner warped regions (Marino et al., 2015; Casassus et al., 2018). Furthermore, warps have even been invoked to explain the variability of active galactic nuclei (AGN) (Raj & Nixon, 2021; Kaaz et al., 2023) and the precession of jets (Miller-Jones et al., 2019; Cui et al., 2023). Tantalizing direct evidence for warps has also been provided by the warped maser emission in NGC 4258 (Miyoshi et al., 1995) and the distorted midplane of the protostellar disks L1489 IRS (Sai et al., 2020) and IRAS 04368+2557 (Sakai et al., 2019).
@@
This prevalence of warped systems demands a rigorous theoretical understanding to model their evolution in space and time. Whilst the early studies laid much of the geometrical framework for global warp evolution (see Petterson, 1977a, b, 1978), they assumed that the disk behaves as a series of rigid, concentric annuli that interact in a purely viscous fashion. The paradigm was fundamentally changed by Papaloizou & Pringle (1983), who realized that the distorted geometry actually establishes pressure gradients which drive internal, radial oscillatory shear flows that are odd about the midplane and efficiently communicate angular momentum. The pressure forces experienced by a fluid parcel vary over the orbital timescale as it traverses the warped geometry. This variation yields a near-resonant forcing, which matches the natural epicylic frequency. To regulate the resonant response, analytical work typically invokes the viscous $\alpha$ prescription (Shakura & Sunyaev, 1973) as a source of damping. Indeed, provided $\alpha$ exceeds the aspect ratio of the disk $H/R$, the viscosity effectively moderates the internal flows and sets up a quasi-steady ‘sloshing’ state whilst the warp evolves over longer timescales – this is the so-called diffusive regime. In the opposite case that $\alpha<H/R$, the weaker viscosity is unable to temper the internal flows and instead the warp is detuned from resonance by dynamically evolving as a bending wave in the wavelike regime (Papaloizou & Lin, 1995). These linear theories were subsequently extended toward nonlinear warp amplitudes by Ogilvie (1999) for the diffusive regime and by Ogilvie (2006, 2018) for the wavelike regime. Efforts to combine both regimes into a generalized set of evolutionary equations have also recently been pursued (Martin et al., 2019; Dullemond et al., 2022).
@@
All of these theories rely on the simplification that the disk flow structure is laminar and smooth and that any small-scale turbulence is well parameterized by the viscous $\alpha$ model. With this in mind, a number of global, laminar numerical experiments have found good agreement with the theoretical equations for warp evolution in both regimes (e.g. Lodato & Price, 2010; Fragner & Nelson, 2010; Facchini et al., 2013; Nealon et al., 2015; Kimmig & Dullemond, 2024; Fairbairn & Ogilvie, 2023). However, such a simple viscous closure model might not capture the true behavior of the underlying turbulence and its interplay with the warp. Indeed, it has long been appreciated that distorted disks are susceptible to a hydrodynamic, parametric instability which feeds off the internal, oscillatory shear flows (Papaloizou & Terquem, 1995; Gammie et al., 2000). This small-scale instability is most readily understood within a local warped shearing box model (Ogilvie & Latter, 2013a, b) which was numerically implemented in the two-dimensional simulations of Paardekooper & Ogilvie (2019). They recovered the linear growth phase of the instability and the ensuing nonlinear saturated state. This significantly modified the laminar flows and reduced the associated angular momentum transport in the warped disk. Furthermore, they found complicated variability in the final turbulent state, as a function of the warp amplitude, and also evidence of hysteresis as the saturated flows depended on the warp history. This instability can also efficiently extract energy from the sloshing flows and lead to a rapid damping of the driving warp (Fairbairn & Ogilvie, 2023). Unfortunately, this richness of dynamics is typically unresolved in global numerical experiments. The only clear identification so far has been the high-resolution, 120 million Lagrangian particle runs of Deng et al. (2020) and Deng & Ogilvie (2022), which attains $\sim 8$ elements per scale height. The emergence of the parametric instability dramatically damps the warp, strongly departing from the laminar prescriptions.
@@
This purely hydrodynamic picture is further complicated by the possible presence of magnetic fields in warped disks – for example, relevant to the case of X-ray binaries or AGN disks. Analytical progress on this front is limited to the regime of strong magnetization for which vertical field lines threading the disk provide an effective tension – resisting the oscillatory, shearing flows and detuning the aforementioned Keplerian resonance (Paris & Ogilvie, 2018). In the alternate case of weak fields, the disk is readily unstable to the magnetorotational instability (MRI) (Balbus & Hawley, 1991; Hawley & Balbus, 1991). Whilst this mechanism is often invoked to explain radial transport of mass and angular momentum, the turbulence might also help moderate the vertically shearing, sloshing flows. Indeed, often the in-plane accretion stresses are equated to some value of effective $\alpha$, which is then used in the isotropic, viscous theories outlined above. In reality, the picture is inevitably more complicated as the MRI turbulence is anisotropic and perhaps does not respond instantaneously to the local shear rate. Therefore, we need to appeal to numerical simulations to gain further insight. Indeed, the global, magnetized warped torus simulations of Sorathia et al. (2013a) suggested a distinct evolution from that predicted by viscous theory. Furthermore, they found that the flow was dominated by hydrodynamic rather than magnetized stresses, akin to their purely hydrodynamic runs in Sorathia et al. (2013b). Contrastingly, Nealon et al. (2016) were able to reproduce the MHD simulations of a warped disk around a spinning black hole (Krolik & Hawley, 2015) using viscous simulations. Despite these global efforts, there has been little work to investigate in detail how the local internal flows are modified by the MRI. Exceptionally, Torkelsson et al. (2000) performed three-dimensional stratified, shearing box simulations of an MRI turbulent disk where the effects of the warp are mocked up via an imposed epicylic, sloshing initial condition that is odd about the midplane. This controlled experiment identified the early emergence of the parametric instability, followed by exponential damping of the sloshing motions due primarily to the hydrodynamic transport of momentum. The damping timescales found are consistent with vertical viscosities which are (coincidentally) comparable to the in-plane MRI effective viscosity – thus supporting an isotropic $\alpha$.
@@
Despite this progress, there are several routes demanding further inquiry. The MRI simulations of Torkelsson et al. (2000) implement the warp as a free initial condition that quickly decays. Therefore, there is little time to establish a saturated, magnetized turbulent state in which the parametric instability, MRI, and sloshing flows come into an equilibrium. Indeed, extracting the nature of the turbulent stresses is complicated by the limited timescales. Contrastingly, the detailed parametric instability study of Paardekooper & Ogilvie (2019) establish steady sloshing states but focus on two-dimensional hydrodynamics. In this paper, we will find a compromise between the two approaches. In particular, we will compare controlled hydrodynamic and magnetized, local simulations of oscillatory shear flows that are maintained by a time-dependent forcing function, mimicking the radial pressure gradients set up by warped disks. In this initial study, we will further simplify matters by considering the unstratified problem, which will be extended toward the stratified case in future work. The outline for this paper is as follows. In Section 2 we will describe the problem setup and summarize the laminar model predictions. In Section 3, we will benchmark our numerical setup by capturing the linear growth of the parametric instability and compare it with theoretical predictions. In Section 4, we will examine the fiducial hydrodynamic sloshing run and study the emerging stresses. This will be compared with the fiducial magnetized sloshing run in Section 5, before we perform a study varying the forcing amplitude. We will discuss our findings in the context of previous studies in Section 6 and finally conclude in Section 7.
@@ 2 Problem Setup
where $\rho$ is the density, $\mathbf{u}$ is the velocity vector, and $\mathbf{B}$ is the magnetic field. The total pressure $P^{*}=p+\mathbf{B}\cdot\mathbf{B}/2$, is composed of the magnetic pressure and the thermal pressure $p$. Throughout this work, we will assume the simplest isothermal equation of state (EoS) $p=c^{2}\rho$, connecting pressure and density via the sound speed $c$. This removes the need to explicitly solve an energy equation. On the right-hand side of equation (2), we see contributions from the Coriolis force and the tidal potential, featuring the shear parameter $q=-d\ln\Omega/d\ln r=3/2$, for a Keplerian disk. Furthermore, the shear viscous stress tensor (Landau & Lifshitz, 1959) is defined as
@@
$\mathsf{T}_{ij}=\rho\nu_{ij}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}-\frac{2}{3}\delta_{ij}\nabla\cdot\mathbf{u}\right),$ (4)
@@
where we allow for an anisotropic kinematic viscosity $\nu_{ij}$. [^1] Note that symmetry imposes $\nu_{ij}=\nu_{ji}$. Furthermore, this simple prescription for characterizing anisotropy means that $\mathsf{T}_{ij}$ does not generally transform in a consistent tensorial fashion. Another nonideal effect is included in the induction equation (3) when the ohmic resistivity $\eta$ is set to be nonzero. Note that for the purposes of this focused study, we have neglected the vertical component of gravity and therefore consider unstratified disk setups. This extension will be addressed in future work.
@@
Figure 1: Global and local cartoon of a warped disk. Left: The global warped disk consists of a series of orbits with a radially varying orientation of angular momentum. Right: A zoomed-in patch of the warped disk manifests as a tilting of the midplane. This projects a component of the vertical pressure gradient in the radial direction and drives sloshing flows.
@@
This study is primarily interested in probing the effect of instabilities and ensuing nonlinear turbulence in moderating oscillatory flows forced by distorted geometries. In particular, it has long been appreciated that warped disk evolution is driven by self-induced internal flows (e.g. Papaloizou & Pringle, 1983; Ogilvie & Latter, 2013a). To gain some heuristic intuition, consider the warped disk cartoon in Fig. 1. The globally distorted geometry is visualized on the left, where the midplane changes inclination as a function of the radius. The orientation of the angular momentum on each disk annulus is formally encoded by the radial and temporal variation of the unit tilt vector $\hat{\mathbf{l}}(r,t)$. The warp amplitude is then quantified by $\psi=\left|d\hat{\mathbf{l}}/d\ln r\right|$. On the right, we zoom in to a portion of the disk and examine the local manifestation of the warp as a tilted midplane, defined by $z_{\rm mid}=\psi x_{\rm mid}$. In a flat, stratified disk, the vertical hydrostatic pressure gradient is given by $\partial p/\partial z=-\rho\Omega_{0}^{2}z$. Therefore, upon tilting the midplane, a portion of this force is projected into the radial direction, yielding $\mathbf{\hat{x}}\cdot\nabla p\sim\rho\Omega_{0}^{2}\psi z$ (Lodato & Pringle, 2007). As we track the orbit around the disk, the midplane tilts back and forth, incurring a sinusoidal dependence in the radial forcing such that it has the form $f_{\rm w}\sim\rho\Omega_{0}^{2}\psi z\cos(\Omega_{0}t)$. The characteristic properties of this forcing are that it is linear in the vertical coordinate $z$ and oscillates on the orbital timescale.
@@
Of course, the simple, unstratified shearing box setup introduced above precludes the self-consistent generation of these pressure gradients. Instead, we will implement an artificial forcing function that emulates key properties in a controlled fashion. To this end, we will add
@@
to the right-hand side of equation (2), where $\psi_{\rm f}$ is now a complex number allowing for an arbitrary choice of phase. Here, $L_{z}$ is the vertical domain in our shearing box, whilst $H\equiv c/\Omega_{0}$ sets the characteristic scale height. Near the box midplane, this form approximates a linear function of $z$ before turning around to satisfy periodic boundary conditions. This idealized model, although simple, contains all of the fundamental physics for driving sloshing flows in warped disks. An alternative approach might be to use the warped generalization of the shearing box, developed by Ogilvie & Latter (2013a), which explicitly incorporates the warped geometry via Lagrangian coordinates that track the tilted orbital motion. However, in order to leverage current state-of-the-art MHD numerical code bases, with inbuilt support for standard shearing boxes, we will adopt our simplified approach.
To gain some intuition about the growth of sloshing motions in our setup, consider the hydrodynamic limit of equations (1) and (2), equipped with the forcing function in equation (5). Furthermore, we will assume that the flow structure is horizontally homogeneous, laminar, and moderated by the shear viscosity, which acts as a model for the turbulent stresses investigated later in this paper. Motivated by the sinusoidal vertical structure of $\mathbf{f}_{\rm w}$ the velocity response follows
@@
where $u_{\rm K}=-q\Omega_{0}x$ is the usual Keplerian shear flow solution in the absence of warped forcing, and we have defined $k_{\rm b}\equiv 2\pi/L_{z}$. It is also sometimes useful to work with the velocity relative to the background Keplerian flow, which we will denote $\mathbf{v}=\mathbf{u}-\mathbf{u}_{\rm k}$. Inserting this ansatz into the governing equations yields
@@
In the simplest case for which the system is unforced and inviscid, the dynamics reduces to the epicylic oscillator solution
@@
for which $S$ sets the sloshing amplitude and phase. Note that equation (11) represents an exact solution to the (unforced) nonlinear shearing box equations, absent magnetic fields. The vertical layers oscillate independently as a series of stacked epicylic orbits. The amplitude of these oscillations as a function of $z$ is enforced by our ansatz in equations (6) and (7), although for this free sloshing, any arbitrary function of $z$ is obviously permitted. We will restrict ourselves to this coherent sinusoidal structure, which is established by our forcing function. A useful diagnostic which encodes the magnitude of this xxyy sloshing motion is the epicylic amplitude
@@
When the system is inviscid but now forced, equation (2) reduces to the standard, resonantly forced oscillator with a linearly growing response
@@
This resonance between the warp-induced radial forcing frequency and the orbital frequency (equal to the natural epicyclic frequency for our Keplerian shear) underlies the difficulty in developing a theory for warps in the low viscosity, bending wave regime, where the response is dynamical and no quasi-steady sloshing state can be found. Meanwhile, the inclusion of viscosity tempers the resonance and gives the steady-state sloshing
@@
$B_{\rm vis}(t)=-\frac{\Omega_{0}^{2}\psi_{\rm f}}{2\left[\nu_{xz}\nu_{yz}k_{\rm b}^{4}+i\Omega_{0}(\nu_{xz}+\nu_{yz})k_{\rm b}^{2}\right]}e^{i\Omega_{0}t}.$ (16)